4. The mean hemoglobin level in the blood for a certain large group of individuals is 21.0 grams per milliliter (g/ml). The standard deviation is 2 g/ml. If a sample of 25 individuals is selected, find the probability that the mean will be greater than 21.3 g/ml. Assume that the variable is normally distributed.

1
Expert's answer
2022-03-20T06:41:24-0400

We have a normal distribution, "\mu=21.0, \sigma=2, n=25."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}" =

"=\cfrac{21.3-21.0}{2/\sqrt{25}}=0.75,"

"P(\bar{X}>21.3)=1-P(\bar{X}<21.3)=\\
=1-P(Z<0.75)="

"=1-0.7734=0.2266" ​(from z-table)